OpenStudy (anonymous):

What is the greatest common factor of 19x7 and 3x5

OpenStudy (anonymous):

19x 7 and 3x 5 GCF = x^5 because: 19x^7 = x^5(19x^2) 3x^5 = x^5(3

OpenStudy (anonymous):

:]